# Learning Priors for Non Rigid SfM from Casual Videos

### v1

@@ Abstract
We tackle the long-standing challenge of reconstructing 3D structures and camera positions from videos. The problem is particularly hard when objects are transformed in a non-rigid way. Current approaches to this problem make unrealistic assumptions or require a long optimization time. We present TracksTo4D, a novel deep learning-based approach that enables inferring 3D structure and camera positions from dynamic content originating from in-the-wild videos using a single feed-forward pass on a sparse point track matrix. To achieve this, we leverage recent advances in 2D point tracking and design an equivariant neural architecture tailored for directly processing 2D point tracks by leveraging their symmetries. TracksTo4D is trained on a dataset of in-the-wild videos utilizing only the 2D point tracks extracted from the videos, without any 3D supervision. Our experiments demonstrate that TracksTo4D generalizes well to unseen videos of unseen semantic categories at inference time, producing equivalent results to state-of-the-art methods while significantly reducing the runtime compared to other baselines. Project page: https://tracks-to-4d.github.io.
Keywords:
Structure From Motion Dynamic Videos 3D Reconstruction Equivariance Symmetries
Figure 1: We present TracksTo4D, a method for mapping a set of 2D point tracks extracted from casual dynamic videos into their corresponding 3D locations and camera motion. At inference time, our network predicts the non-rigid structure and camera motion in a single feed-forward pass. Top. An example showing the input video frames ordered from left to right. A set of pre-extracted point tracks that are used as input to our network and presented in corresponding colors. Bottom. The outputs of our network for 3 time steps as observed from the right side of the scene. The camera trajectory is present as gray frustums, whereas the current camera is marked in red. The reconstructed 3D scene points are presented in corresponding colors to the input tracks on the top. See video visualizations here.
@@ 1 Introduction
Predicting 3D structures in dynamic scenes is a challenging problem. In contrast to rigid scenes where the epipolar geometry constraints hold between the corresponding points of different views [17], determining the depth of a moving point from monocular views is an ill-posed problem. Moreover, while multiview camera pose estimation can be solved successfully by Structure from Motion methods [37, 53, 30], these techniques tend to fail in scenarios with significant non-rigid motion [21].
Background and previous work.
Unfortunately, current approaches to solving this problem are based on unrealistic assumptions or suffer from long optimization times per scene that make them inapplicable to real-world problems. More concretely, early techniques for Non-Rigid SFM [6] used matrix factorization of long-term point tracks matrix of the non-rigid scene part. This was done by assuming that a small linear basis and per-frame coefficients can formulate the dynamic structure. By further assuming an orthographic camera is used, they omitted camera translations and solved for rotations and 3D structure by matrix factorization. While follow-up work improved accuracies [22, 11], this setup is still not practical for real-world casual videos, captured by pinhole cameras. Recently, the Neural Radiance Field (NeRF) [28] concept has brought major progress to novel view synthesis and surface reconstruction [60, 49] for static scenes. While some follow-up work extended NeRF to non-rigid scenes [23, 24, 33], optimizing a non-rigid NeRF requires a long, and per-scene optimization. Other recent approaches [26, 21, 62, 63] optimized for consistent depth maps given a specific monocular video by fine-tuning pre-trained single-view-depth-predictions [34]. These methods showed impressive progress in camera and non-rigid structure estimation, but also require per-video optimization at inference time. Other approaches learned prior and parameterizations for specific non-rigid object classes such as humans [36, 25]. This prior enables shorter inference times for obtaining 3D structures from monocular videos that contain instances from the learned category [55]. However, this approach does not generalize well across shape categories and mostly requires per-video finetuning.
@@
In this work, we propose TracksTo4D, a novel deep learning-based approach for predicting the 3D dynamic Non-Rigid Structure And camera Motion from videos. Our key observation is that the 2D motion present in casual videos shares common patterns across a variety of semantic video categories. Crucially, and as was suggested by [6], this motion can be effectively encoded into a point track matrix, where each entry represents the 2D position of a tracked point in a specific video frame. Rather than processing image pixels or semantic features that vary dramatically across different scenes and objects, we propose directly processing these point track matrices to learn generic, class-agnostic features. This also differs from the common approach of using optical flow representations of motion, as utilizing optical flow fields as input to deep neural networks is non-trivial.
@@
TracksTo4D  builds on the recent progress in point tracking [12, 18], which introduced the ability to extract long-term point tracks from casual videos. We leverage this progress and design a deep neural network architecture, that maps the point tracks of a given video into corresponding camera poses, 3D points, and per-point non-rigid motion level values (See Fig. 1). Crucially, the mapping we aim to learn preserves two natural symmetries: (1) the points being tracked can be arbitrarily permuted without affecting the problem; (2) the frames containing these points exhibit temporal structure, adhering to an approximate time-translation symmetry. Following the Geometric Deep Learning paradigm [7], we build upon recent theoretical advances in equivariant learning [27] and integrate these two symmetries into our network architecture using dedicated attention mechanisms.
Predicting the camera poses and the 3D non-rigid motion from 2D motion alone is an ill-posed problem without any constraints, and equivalent to finding the 3D structure from each frame [3]. Inspired by previous work [6], we model the predicted 3D point cloud as a weighted sum of a 3D structure basis, where the first basis is assumed to fully represent the rigid points, and the rest of the basis elements are used to complete the non-rigid deviation for the non-rigid points. The predicted motion level values determine the non-rigidity of each 3D world point, i.e. the inconsistency of the rigid assumption that is represented by the first basis element for this world point.
@@
In our experiments, TracksTo4D is trained on the Common Pets dataset [38] using only 2D point tracks extracted by [18], without any 3D supervision by simply minimizing the reprojection errors. Interestingly, we observe that by just "watching" point tracks in videos, our network learns to infer their 3D location and the camera motion. We evaluate our method on test data with GT depth maps and demonstrate that it generalizes well, and produces comparable results to the state-of-the-art methods, with significantly shorter inference time.
@@
In summary, our contributions are: (1) Designing an equivariant network architecture that maps from 2D point trajectories into corresponding non-rigid structures and camera poses; (2) Introducing a training paradigm to identify the rigid part of the scene for constraining the camera poses; (3) Simple training on casual videos without any 3D supervision; (4) Demonstrating fast inference time, accurate results, and strong generalization across animal categories.
2 Related Work
Structure from Motion
The Structure from Motion (SfM) problem recovers the 3D structure and the camera poses from images. Early methods [44, 39] factorized a 2D measurement matrix of corresponding image points into a multiplication of a camera poses matrix by a 3D points matrix. These methods assume that each recovered 3D point is tracked over the entire image sequence which is not the case in practical use cases. Incremental approaches for SfM [37, 1, 53] start from a minimal set of images, to build an initial reconstruction. Then, the rest of the images are added to the scene one by one, by finding 2D correspondences from the new image into scene points. To prevent drift, Bundle Adjustment [45] is repeatedly applied to the reconstruction during this process. Several global methods [52, 10, 19, 51, 32] reduced the running time by using relative two-view geometries between image pairs to extract absolute camera poses in a single step, followed by a final round of Bundle Adjustment.
Recently, Moran et al. [29] presented a method for applying an order equivariant architecture for mapping a set of multiview image correspondences into two sets of cameras and 3D points. It was trained on a dataset of scenes and obtained a reduced optimization time on novel scenes, due to the learned prior. A follow-up work [8] improved generalization by adding graph attention layers into the architecture of Moran et al. Other recent deep learning-based techniques used differentiable Bundle Adjustment layers [47, 40], or diffusion models [61, 48]. The above-mentioned techniques assumed that the structure is rigid such that the whole scene can be represented in a single 3D structure.
Simultaneous Localization and Mapping (SLAM)
In contrast to SfM which assumes unordered images, monocular Simultaneous Localization and Mapping (SLAM) methods [30, 31, 15] get as input a sequence of video frames recorded by a single camera. These methods often define a set of keyframes for reconstructing the scene with a mapping thread, while registering every new frame into the structure by a localization thread. The mapping thread repeatedly applies Bundle Adjustment for updating the structure. ORB-SLAM [30] uses feature points for mapping a sparse scene representation and tracks the camera with sparse point matches. Other methods [31, 15, 14] use denser scene representations and align cameras directly by using photometric constraints.
More recent methods combine classical approaches with deep learning-based tools e.g. depth prediction from a single image, optical flow prediction, or relative camera pose estimation [59, 5, 50, 65, 66, 41]. As in SfM, these methods assume rigid scenes, fail to produce the cameras in scenes with large portions of non-rigid motion, and cannot reproduce non-rigid parts of the scene. DROID-SLAM [43] used synthetic data with ground truth 3D supervision for learning to predict camera poses via deep-based bundle adjustment on keyframes while excluding dynamic objects. ParticleSfM [64] pre-trained a motion prediction network on a synthetic dataset with ground truth supervision, and used it to filter out any non-rigid content for reproducing the cameras in dynamic scenes. Both, [43, 64] do not infer the non-rigid 3D structure.
Test-Time Optimization for Non-Rigid Scenes
Recent methods [26, 21, 62, 63] finetuned the monocular depth estimation from a pre-trained model [35, 34] using optical flow constraints [42], for obtaining consistent dense depth maps for a monocular video. [62] further optimized motion maps that represent the $\gamma$ parameters of zero mean Cauchy distributions for handling scenes with large dynamic motion. [54, 16] use depth from single image estimations, to improve novel view synthesis in dynamic scenes. [24] further optimized for the unknown camera poses together with the dynamic radiance field optimization. LASR [56], ViSER [57] and BANMo [58] optimize for a dynamic surface given a monocular video by assuming rigid bones and linear blend skinning weights. However, all the above-mentioned methods require per-scene optimization, resulting in slow inference. Recently, [38] presented the Common Pets in 3D (COP3D) dataset that contains casual, in-the-wild videos of pets, and used it to learn priors for novel view synthesis in dynamic scenes.
3 Problem formulation
Our problem setup is illustrated in Fig. 2. Given a video of $N$ frames, let $M\in\mathbb{R}^{N\times P\times 3}$ be a pre-extracted point tracks tensor (Fig. 2, left side). This tensor represents the two-dimensional information about a set of $P$ world points that are tracked throughout the video. Each element in the tensor, $M_{i,j}$, stores three values: $(x,y,o)$ where $x,y\in\mathbb{R}$ are respectively the observed horizontal and vertical locations of point $j$ in frame $i$, and $o\in\{0,1\}$ indicates whether point $j$ is observed in frame $i$ or not. Our goal is to learn a function $f$ to map the input point tracks tensor $M$ into a set of per-frame camera poses $\{R_{i},\mathbf{t}_{i}\}_{i=1}^{N}$ and per-frame 3D points $\{X_{i}\}_{i=1}^{N}$, where $R_{i}\in SO(3),\mathbf{t}_{i}\in\mathbb{R}^{3},X_{i}\in\mathbb{R}^{P\times 3}$ (Fig. 2, right side).
Discussion.
We note that given enough point tracks, $M$ contains the 2D motion constraints that can be provided by the video. The source of each 2D motion constraint can be either from a global camera 3D motion or from an independent object motion. While the pixel appearance and semantic information could be used as well as inputs in our task, we find that the 2D motion signal $M$ gives useful geometric constraints for our setup. Moreover, this abstraction of the video helps our network generalize well to out-of-distribution videos with object classes that have not been observed during training but share similar 2D motion characteristics as we show in our experiments. Therefore $f$ only uses $M$ as input and does not rely on the input video once $M$ has been extracted.
Figure 2: Our pipeline Our network takes as input a set of 2D point tracks (left) and uses several multi-head attention layers while alternating between the time dimension and the track dimension (middle). The network predicts cameras, per-frame 3D points, and per-world point movement value (right). The 3D point colors illustrate the movement level values, such that points with high/low motion values are presented in blue/purple colors respectively. These outputs are used to reproject the predicted points into the frames for calculating the reprojection error losses. See details in the text.
@@ 4 Method
4.1 Overview
Our pipeline is illustrated in Fig. 2. Our method receives $M\in\mathbb{R}^{N\times P\times 3}$ as input. This tensor is being processed by a neural architecture composed of multi-head attention layers where the attention is calculated sequentially on the $P$ and the $N$ dimensions in each layer (Sec. 4.3). After several such layers, we perform a projection to $N$ per-frame parameters and $P$ per-world-point parameters which determine the camera poses and 3D dynamic structure (Sec. 4.2). This projection is done by first pooling over the secondary dimension and applying suitable layers to each one of the outputs to get the per-camera and per-world points parameters respectively. Our network is trained by minimizing the reprojection error and other regularization losses (Sec. 4.4) that are used to update the model parameters.
4.2 Constraining the Non-Rigid Structure and Motion
Given the input 2D tracks, we wish to describe their motion by global camera motion and non-rigid self-motion. The 2D motion that is produced from rigid points constrains well the camera motion, but predicting the camera poses and the 3D non-rigid motion from 2D motion alone is an ill-posed problem without any constraints [3].
We tackle this challenge in two ways. First, as done in classical Non-Rigid Structure from Motion [6], we constrain the output points to be formulated by a linear combination of basis elements. This constrains the 3D motion to be simple and prevents "jumps" in the unconstrained depth. More formally we assume that the non-rigid structure is formulated by a few basis elements $B_{1},\dots,B_{K}\in\mathbb{R}^{P\times 3}$ such that $X_{i}=\sum_{k=1}^{K}c_{ik}B_{k}$ where $\{c_{ik}\}_{k=1}^{K}$ are the $K$ coefficients which define the 3D structure of frame $i$.
Secondly, we define the first basis element $B_{1}\in\mathbb{R}^{P\times 3}$ to be a rigid approximation of the points, such that the first basis coefficient in every frame is $1$, i.e. $c_{i1}=1$ for $i=1,\dots,N$. This helps in constraining the camera motion from the 2D motion of the rigid scene part. However, since this global rigid approximation cannot explain the 2D motion that is produced by the non-rigid world points, we model the reprojection error of the rigid approximation with $P$ zero-mean Cauchy distributions. For a non-rigid world point, $B_{1}$ by itself cannot produce low reprojection errors, and therefore its associated $\gamma$ is expected to be higher. The parameters of the Cauchy distributions, $\gamma_{1},\dots\gamma_{P}\in\mathbb{R}_{+}$ (Fig. 2, "Movement Level Values") are predicted as well by the network and our training maximizes the likelihood of this assumption (see details in Sec. 4.4). This allows our network to benefit from the rigid part for camera pose prediction while still being robust to the non-rigid motion. The rest of the basis elements determine the non-rigid deviation of the non-rigid points from the rigid approximation.
In summary, our function $f$ gets $M\in\mathbb{R}^{N\times P\times 3}$ as an input and predicts per-frame outputs:
$R_{i}\in SO(3),\mathbf{t}_{i}\in\mathbb{R}^{3},\{c_{ik}\}_{k=2}^{K},\text{for}\ \ i=1,\dots,N$ (1)
and per-point outputs:
$B_{1},\dots,B_{K}\in\mathbb{R}^{P\times 3},\gamma_{1},\dots,\gamma_{P}\in\mathbb{R}_{+}$ (2)
which together define the non-rigid structure and motion.
4.3 Architecture
@@
Our input is a tensor $M\in\mathbb{R}^{N\times P\times 3}$ representing a sequence of $N$ frames each with $P$ point coordinates. This structure gives rise to two key symmetries: First, the order of the $P$ points within each frame does not matter - permuting this axis results in an equivalent problem [27]. Formally, this axis has a permutation symmetry $S_{P}$ where $S_{P}$ is the symmetric group on $P$ elements. Second, along the temporal $N$ axis, we have an approximate translation symmetry arising from the ordered video sequence. This means that translating the time frames is required to result in nearly the same overall representation. We model this with a cyclic group $C_{N}$ of order $N$. Both symmetries are illustrated in Fig. 3.
@@
Taken together, the full symmetry group of the input space is the direct product $\mathcal{G}=C_{N}\times S_{P}$ combining these frame and point permutation symmetries  [^1] This is different from the symmetry group studied in [29], where the temporal structure was not exploited. Next, we will design an architecture equivariant to $\mathcal{G}$, so that we can incorporate these domain priors and ensure that the model takes into account the underlying data structure.
Network Architecture.
Building on the analysis from the Deep Sets for Symmetric elements (DSS) approach [27], if we assumed we can use only linear layers, our architecture would take the form:

### v2

# Fast Encoder-Based 3D from Casual Videos via Point Track Processing

@@ Abstract
This paper addresses the long-standing challenge of reconstructing 3D structures from videos with dynamic content. Current approaches to this problem were not designed to operate on casual videos recorded by standard cameras or require a long optimization time. Aiming to significantly improve the efficiency of previous approaches, we present TracksTo4D, a learning-based approach that enables inferring 3D structure and camera positions from dynamic content originating from casual videos using a single efficient feed-forward pass. To achieve this, we propose operating directly over 2D point tracks as input and designing an architecture tailored for processing 2D point tracks. Our proposed architecture is designed with two key principles in mind: (1) it takes into account the inherent symmetries present in the input point tracks data, and (2) it assumes that the movement patterns can be effectively represented using a low-rank approximation.TracksTo4D is trained in an unsupervised way on a dataset of casual videos utilizing only the 2D point tracks extracted from the videos, without any 3D supervision. Our experiments show that TracksTo4D can reconstruct a temporal point cloud and camera positions of the underlying video with accuracy comparable to state-of-the-art methods, while drastically reducing runtime by up to 95%. We further show that TracksTo4D  generalizes well to unseen videos of unseen semantic categories at inference time. Project page: https://tracks-to-4d.github.io.
@@ 1 Introduction
Predicting 3D geometry in dynamic scenes is a challenging problem. In this problem setup, we are given access to multiple images of a scene taken sequentially, e.g., from a monocular video camera, where both the content in the scene and the camera are moving. Our task is to reconstruct the dynamic 3D positions of the points seen in the images and the camera poses. This fundamental problem has gained significant interest from the research community over the years [6, 34, 23, 59], mainly due to its important applications in many fields such as robot navigation, autonomous driving and 3D reconstruction of general environments [17]. Importantly, in contrast to static scenes where the epipolar geometry constraints hold between the corresponding points of different views [15], determining the depth of a moving point from monocular views is an ill-posed problem [3]. This causes standard Structure from Motion techniques [38, 51, 30] to be inadequate in this setup [22].
Previous work and limitations.
Many existing approaches for the above problem make simplifying assumptions that limit their applicability to real-world scenarios. Methods based on orthographic camera models and low-rank assumptions use matrix factorization techniques [6, 23], but the orthographic camera assumption might not be realistic and may cause reconstruction errors. Techniques that incorporate depth priors often require lengthy optimization processes in order to make the depth estimates across frames consistent [22, 59]. Other physics-based approaches make assumptions about rigid bones [56, 54] or isometric deformable surfaces [34] and typically involve complex, slow optimization per video. In addition, they may require foreground-background segmentation of the moving content, which is not always easily obtained. Alternatively, some methods are specifically tailored to certain object classes like humans [48], restricting their domain to those limited cases. Consequently, these prior methods are either not directly applicable to casual videos, or require long optimization time per video.
Figure 1: We present TracksTo4D, a method for mapping a set of 2D point tracks extracted from casual dynamic videos into their corresponding 3D locations and camera motion. At inference time, our network predicts the dynamic structure and camera motion in a single feed-forward pass. Our network takes as input a set of 2D point tracks (left) and uses several multi-head attention layers while alternating between the time dimension and the track dimension (middle). The network predicts cameras, per-frame 3D points, and per-world point movement value (right). The 3D point internal colors illustrate the predicted 3D movement level values, such that points with high/low 3D motion are presented in red/purple colors respectively. These outputs are used to reproject the predicted points into the frames for calculating the reprojection error losses. See details in the text. The reader is encouraged to watch the supplementary video visualizations.
@@
We propose TracksTo4D,1114D since we have three Euclidean coordinates with an additional time coordinate a novel approach for fast reconstruction of sparse dynamic 3D point clouds and camera poses from casual videos. Our main idea is to train a neural network on multiple videos to learn the mapping from the input image sequence to a sequence of the scene’s 3D point clouds and camera poses. After training, the trained network can be efficiently applied to new image sequences using a single feed-forward pass, avoiding costly optimization.
To enhance the method’s ability to generalize across different types of videos and scenes, we made a crucial design choice: our approach processes point track tensors as input, rather than operating directly on the image sequence. Specifically, Each entry $(n,p)$ in these tensors represents the 2D position of a tracked point $p$ in a specific video frame $n$ [6]. Our main insight is that point track tensors may exhibit more common motion patterns across casual video domains compared to image pixels. In other words, we argue that processing the raw point track data rather than scene-specific pixels or features may enable learning class and scene-agnostic internal feature representations for improved generalization. Importantly, recent advances in point tracking [10, 18] enable efficiently inferring these point tracks from casual videos using pre-trained models. These two properties make point track matrices an attractive input for our learning method.
Following this design choice, we design our architecture according to two principles: (1) process point track tensors, which have a unique structure, and (2) encode meaningful prior knowledge about the reconstruction problem, as the problem is ill-posed in general. In the following, we address these desired properties.
@@
First, we design a network architecture that can effectively and efficiently handle point track inputs. To do that, we propose a novel layer design that takes into account the symmetries of the problem: the mapping we aim to learn, from point track matrices to 3D point clouds and camera poses, preserves two natural symmetries: (i) the points being tracked can be arbitrarily permuted without affecting the problem; (ii) the frames containing these points exhibit temporal structure, adhering to an approximate time-translation symmetry. Following the Geometric Deep Learning paradigm [7], we build upon recent theoretical advances in equivariant learning [27] and integrate these two symmetries into our network architecture using dedicated attention and positional encoding mechanisms.
Second, a key challenge in predicting 3D dynamic motion and camera poses from 2D point tracks is that this problem is inherently ill-posed without additional constraints [3]. To address this, we integrate a low-rank movement assumption into our architecture, following the seminal work of [6] which constrained output point clouds to be linear combinations of basis elements. Specifically, given an input point track tensor, our architecture equivariantly predicts a small set of input-specific basis elements. The output point clouds at each time frame are then defined as a linear combination of these basis elements, with the coefficients also predicted by the network. Notably, the first basis is assumed to fully represent the 3D static points in the video, while the remaining basis elements capture the 3D dynamic deviations. This structure effectively restricts the predicted point clouds to have a more specific form, making the problem more constrained.
@@
Our approach is trained on a dataset of extracted point track matrices [18] from raw videos without any 3D supervision by simply minimizing the reprojection errors, aiming to predict output point clouds that, after undergoing a perspective projection, will return the original 2D point tracks. In our experiments, TracksTo4D is trained on the Common Pets dataset [40]. We evaluate our method on test data with GT camera poses and GT depth information for point tracks, and demonstrate that it produces comparable results to state-of-the-art methods, while having a significantly shorter inference time by up to $95\%$. In addition, we show the method’s ability to generalize to out-of-domain videos.
@@
In summary, our contributions are (1) A novel modeling of the dynamic reconstruction problem via learning on point tracks without 3D supervision; (2) A novel deep learning architecture incorporating two key principles: accounting for the symmetry of the data and encoding low-rank structure in the predicted point clouds (3) Experiments demonstrating extremely fast inference time compared to baselines, accurate results, and strong generalization across other categories.
@@ 2 Method
Problem formulation.
Given a video of $N$ frames, let $M\in\mathbb{R}^{N\times P\times 3}$ be a pre-extracted $2D$ point tracks tensor (Fig. 1, left side). This tensor represents the two-dimensional information about a set of $P$ world points that are tracked throughout the video. Each element in the tensor, $M_{i,j,:}$, stores three values: $(x,y,o)$ where $x,y\in\mathbb{R}$ are respectively the observed horizontal and vertical locations of point $j$ in frame $i$, and $o\in\{0,1\}$ indicates whether point $j$ is observed in frame $i$ or not. Our goal is to train a deep neural network to map the input point tracks tensor $M$ into a set of per-frame camera poses $\{R_{i}(M),\mathbf{t}_{i}(M)\}_{i=1}^{N}$ and per-frame 3D points $\{X_{i}(M)\}_{i=1}^{N}$, where $R_{i}(M)\in\mathbb{SO}(3),\mathbf{t}_{i}(M)\in\mathbb{R}^{3},X_{i}(M)\in\mathbb{R}^{P\times 3}$ (Fig. 1, right side).
Overview of our approach.
Our method receives $M\in\mathbb{R}^{N\times P\times 3}$ as input. This tensor is being processed by a neural architecture composed of multi-head attention layers where the attention is applied in an alternating fashion on the $P$ and the $N$ dimensions in each layer. These layers are defined in Sec. 2.1. After a composition of several such layers, the network uses the resulting features in $\mathbb{R}^{N\times P\times d}$ to predict $N$ camera poses in $\mathbb{SO}^{3}\times\mathbb{R}^{3}$ and $N$ point clouds in $\mathbb{R}^{N\times P\times 3}$. These $N$ point clouds are parameterized as a linear combination of $K$ input specific point cloud bases $B_{1}(M),\dots B_{K}(M)\in\mathbb{R}^{P\times 3}$. This is discussed in detail in Sec. 2.2. Our network is trained in an unsupervised way on a dataset of videos by minimizing the reprojection error and other regularization losses (Sec. 2.3) that are used to update the model parameters. Our pipeline is illustrated in Fig. 1
2.1 Equivariant layers for point track tensors
@@
Our input is a tensor $M\in\mathbb{R}^{N\times P\times 3}$ representing a sequence of $N$ frames each with $P$ point coordinates. This structure gives rise to two key symmetries: First, the order of the $P$ points within each frame does not matter - in other words, permuting this axis results in an equivalent problem [27]. Formally, this axis has a permutation symmetry $S_{P}$ where $S_{P}$ is the symmetric group on $P$ elements. Second, along the temporal $N$ axis, we have an approximate translation symmetry arising from the ordered video sequence. This means that shifting the time frames is required to result in the same shift in our output. We model this with a cyclic group $C_{N}$ of order $N$. Both symmetries are illustrated in Fig. 2. We note that while the cyclic group assumption may not be entirely accurate, we still find it useful as it helps us to derive appropriate parametric layers for our data, similar to how the convolutional layer is derived for data with translational symmetries such as images.
@@
Taken together, the full symmetry group of the input space is the direct product $\mathcal{G}=C_{N}\times S_{P}$ combining these time and point permutation symmetries, acting on $\mathbb{R}^{N\times P\times 3}$ by $((t,\sigma)\cdot M)_{n,p,j}=M_{t^{-1}(n),\sigma^{-1}(p),j}$ for $(t,\sigma)\in\mathcal{G}$ [^2] This is different from the symmetry group studied in [29], where the temporal structure was not exploited.. Next, we will design an architecture equivariant to $\mathcal{G}$, to ensure that the model takes into account the symmetries above.
Linear equivariant layers.
Point track tensors can be viewed as a collection of $N$ individual point tracks, each of which exhibits translational symmetry. The scenario where an object comprises a set of elements with their own symmetry group, such as a set of images or graphs, was previously explored in [27]. In that work, the authors characterized the general linear equivariant layer structure in such cases, termed the Deep Sets for Symmetric Elements (DSS) layer. Building on the DSS approach, our basic linear equivariant layer for the point track tensors $M$ would take the form: